OpenStudy (anonymous):

Tina has a cubic container she uses to store all of her pencils. The cubic container has a side length of 8cm. She would like to paint the entire outside of the container a new, shiny blue color. Blue paint comes in cans that contains 10 ounces of paint. Each can of paint costs $4.00. If every ounce of paint covers 2 square cm of the container, how much will Tina have to spend in total to paint the entire outside of the container? PLZZ HELP !!

OpenStudy (welshfella):

first you have to find the area of outside of the cube . That is the area of each side * 6 as there are 6 sides to a cube. can you do that/

OpenStudy (anonymous):

Is it 36 ? @welshfella

OpenStudy (welshfella):

No each side( or face) is a square with side length 8 so the area of one square is 8 * 8 = ? square cms?

OpenStudy (anonymous):

64 cm

OpenStudy (welshfella):

64 square cms - remember its an area and there are 6 of these so total area = 6 * 64 = ? sq cms

OpenStudy (anonymous):

384cm²

OpenStudy (welshfella):

right now to find the cost we have to find how much paint we need to paint the whole cube

OpenStudy (anonymous):

Kk

OpenStudy (welshfella):

2 sq cms af area requires 1 ounce of paint so one can (10 oz) will paint 10*2 = 20 sq cms so divide 384 by use your calculator what do you get?

OpenStudy (anonymous):

I got 18

OpenStudy (welshfella):

* divide 384 by 20

OpenStudy (welshfella):

no its more than that

OpenStudy (anonymous):

Yeah your right its 19.2

OpenStudy (welshfella):

so thats 19.2 cans and cost will be 19.2 * 4 = $76.80

OpenStudy (anonymous):

In part of the question its says the store will not allow her to purchase a decimal of a can

OpenStudy (welshfella):

right - I just thought of that So she'd have to buy 20 cans which is $80.

OpenStudy (anonymous):

Yeah So $80 is my final answer ?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

Alrighty thank you so much @welshfella :) I really appreciate your help :)

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

Alrighty thank you so much @welshfella :) I really appreciate your help :)